Solve the following differential equations:

  1. 8(8-y2)1/2 ln xdx + (8y-5) xdy = 0
  2. dy/dx = ((x2 - y2)/x2))1/2 + y/x
1
Expert's answer
2021-08-12T04:45:46-0400

Solution

1.

Use Method of seperation of variables.

"8(8-y^2)^{\frac12}ln(x)dx" =-(8y-5)xdy

Rewrite as

"8(8-y^2)^{\frac12}lnxdx" =(5-8y)xdy

By separation,

"\frac{8lnx}{x}dx=\frac{5-8y}{(8-y^2)^{\frac12}}dy"

Rewrite as

"\frac{8ln(x)}{x}dx=\frac{5}{\sqrt{8-y^2}}dy-8(\frac{y}{\sqrt{8-y^2}})dy"

Integrate both sides:

"4ln^2(x)+C=5arcsin(\frac{y}{2\sqrt2})+8\sqrt{8-y^2}"

2)

Equation is homogeneous.

Take y=vx

"\frac{dy}{dx}=v+x\frac{dv}{dx}"

By substitution,

"v+x\frac{dv}{dx}=\sqrt{1-v^2}+v"

By simplication,

"x\frac{dv}{dx}=\sqrt{1-v^2}"

Separate by variables:

"\frac{1}{\sqrt{1-v^2}}dv=\frac1xdx"

Integrate both sides:

arcsin(v)=ln(x)+C

But "v=\frac yx" ,Replace back:

"arcsin(\frac yx)=ln(x)+C"